OpenStudy (anonymous):

i will give a medal and become a fan Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 11 feet by 11 feet by 3 inches for a patio if the concrete costs $63.00 per cubic yard? (1 point) $211.75 $70.58 $317.63 $1,905.75 3. The volume of a cylinder is 980 in.3 The height of the cylinder is 20 in. What is the radius of the cylinder? (1 point) 7 in 49 in 327 in 18 in

OpenStudy (anonymous):

i have no idea how many feet are ina yard and inches in a feet. If you tell me I will give you the answer

OpenStudy (anonymous):

1 yd = 3 ft 1 ft = 12 in

OpenStudy (anonymous):

1 foot = 12 inches 1 yard = 36 inches 1 yard = 3 feet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first put everything in yards: 11 feet is 11/3 yards 3 inches is 3/36=1/12 yards so the volume of the slab is: (11/3)*(11/3)*(1/12)=121/108 cubic yards. Since the cubic yard costs 63$, the prize will be (121/108)*63=70.58$

OpenStudy (anonymous):

thank you you made that one look easier then it seems

OpenStudy (anonymous):

for the 2º part: volume of a cylinder is Area*hieght. Height you got = 20 in, volume too =980 cubic in. So: 980=area*20 from here area = 980/20=49 square in area of a circle is = Pi r^2, where r is radius. So 49=Pir^2 from here r^2=49/Pi=15,6 and r= 3.95

OpenStudy (anonymous):

im pretty sure the radius is 7..

OpenStudy (anonymous):

but I get 3.95.....

OpenStudy (anonymous):

980 pie divided by 20 = 49 49 = R2 r=7

OpenStudy (anonymous):

area of a circle is = Pi r^2

OpenStudy (anonymous):

i just noticed the question up their didn't have the pi sign next to 980, but it does on my paper thats why you were calculating pie in the equation and i wasn't, sorry about that

OpenStudy (anonymous):

:). Then it's = 7